solve th esystem of equations using the elimination method.(can you show all steps)
5x+y-z=6
x+y+z=2
3x+y=4

Answer by ankor@dixie-net.com(22740)

solve the system of equations using the elimination method.
5x + y - z = 6
x + y + z = 2
3x + y = 4
:
Add the 1st two equations to eliminates z
5x + y - z = 6
x + y + z = 2
----------------
6x + 2y + 0 = 8
Simplify, divide by 2
3x + y = 4; which is the same as the 3rd equation,
:
Which means there is no solution to this system.